John is 13 years less than 3 times as old as his son. Six years ago, he was 14 years more than twice as old as his son was then. How old is John?
Answer by dabanfield(803)

John is 13 years less than 3 times as old as his son. Six years ago, he was 14 years more than twice as old as his son was then. How old is John?
Let x be the age of John's son and y be John's age. Six years ago John was y-6 year old and his son was x-6. Then we have:
1.) y = 3*x - 13 and
2.) y-6 = 2*(x-6) + 14
Substitute 3x-13 for y (from 1.) in equation 2.):
3x-13 - 6 = 2*(x-6) + 14
3x-19 = 2x-12+14
Solve the above for x and then calculate y = 3x-13.